# On Continuous Optimization for Constraint Satisfaction Problems

### v1

@@ Introduction
$P_{t+1}^{\prime}$ $=P_{t}+\eta\cdot\nabla\mathcal{C}(P_{t})$ (4)
$P_{t+1}$ $=\mathrm{proj}_{\tilde{D}}(P_{t+1}^{\prime})$ (5)
Based on the randomized rounding in Definition 3, we define the projection for Eq. (5) as follows.
Definition 6 (Projection).
The Euclidean projection of $Q\in\mathbb{R}^{|\mathrm{dom}(x_{1})|\times\cdots\times|\mathrm{dom}(x_{n})|}$, onto $P\in\tilde{D}$, is defined as
@@
The projection operation $\mathrm{proj}_{\tilde{D}}(Q)$ can be computed efficiently by separately projecting each $q_{i}$ onto its respective simplex $\tilde{D}_{i}$, leveraging the product structure of $\tilde{D}$. Each sub-projection corresponds to a standard simplex projection problem (Chen and Ye 2011). By substituting Eq. (5) into Eq. (4), we have the gradient mapping:
$g(P)=\frac{1}{\eta}\left(\mathrm{proj}_{\tilde{D}}(P+\eta\cdot\nabla\mathcal{C}(P))-P\right).$
Lemma 3 and Theorem 5 demonstrate that projected gradient ascent is guaranteed to converge to a point where the gradient mapping is small, i.e., $g(P)||_{2}^{2}\leq\epsilon$.
Lemma 3.
Let $\mathcal{C}$ be a multilinear polynomial. Let $n$ be the number of variable, then $N=\sum_{i=1}^{n}|\mathrm{dom}(x_{i})|$. Let $g(\cdot)$ be the projected gradient mapping onto the simplices $\tilde{D}$. Given $y,z\in\tilde{D}$, if the step size satisfies $\eta\leq\frac{1}{|C|\sqrt{N}}$, then the projected gradient step $y=z+\eta{}g(z)$ satisfies:
$\mathcal{C}(y)-\mathcal{C}(z)\geq\frac{\eta}{2}\|g(z)\|^{2}$
Theorem 5 (Convergence Speed).
@@
With the step size $\eta=\frac{1}{|C|\sqrt{N}}$, the projected ascent direction will converge to a $\epsilon$-projected-critical point in $\mathcal{O}\left(\frac{|C|^{2}\sqrt{N}}{\epsilon^{2}}\right)$ iterations.
The projected gradient moves the assignment toward a point associated with a higher function value, and eventually it will converge to a local maximum where the gradient mapping is small. If local maxima were to exist in the interior of the feasible region, i.e., in $\tilde{D}\setminus\delta\tilde{D}$, the quality of the corresponding decoded solution would not be guaranteed; however, we establish the non-existence of such interior local maxima in Lemma 4. Consequently, this result directly implies Theorem 6.
Lemma 4.
Let $\tilde{D}_{\perp}$ be a hyperplane with $\vec{1}$ as the normal vector. For every non-constant $\mathcal{C}$ and a point $P\in\tilde{D}\backslash{\delta\tilde{D}}$, there exist a region size $\epsilon>0$ and two directions, $q^{+},q^{-}\in\tilde{D}_{\perp}$ such that for all $\eta\in(0,\epsilon)$, the following holds.
$\mathcal{C}(P+\eta{}q^{+})>\mathcal{C}(P)$
Theorem 6 (Optimality).
Any interior point $P\in\tilde{D}\backslash{}\delta\tilde{D}$ cannot be a local maximum of $\mathcal{C}(P)$. The maximal points are only attained from the boundary $\delta\tilde{D}$.
To further illustrate these theoretical results, we include a selection of test cases in the supplementary materials, demonstrating how the projected gradient consistently converges to the boundary $\delta\tilde{D}$ across various scenarios.
Implementation Techniques
The primary computational workload of FourierCSP consists of repeatedly executing Algorithm 2 and Algorithm 3 for each constraint during optimization. Since the COP of individual constraints is independent of one another, these computations can be parallelized. Specifically, both algorithms can be compiled into kernel functions and executed concurrently, thereby accelerating the projected gradient step via hardware-level parallelism. Additionally, we leverage autodiff to compute gradients. As a result, it suffices to construct only the top-down evaluation trace. The top-down messages at the terminal nodes of the decision diagram are complementary, i.e., $m_{td}[\texttt{true}]+m_{td}[\texttt{false}]=1$. This allows us to compute the top-down message for only one terminal node while skipping all nodes that do not contribute to its computation. When the COP expression is sufficiently compact, it can be evaluated in closed form without traversing the decision diagram, thereby further reducing computational overhead.
Experimental Results
@@
In this section, we compare our tool, FourierCSP, with systematic and local search solvers. The objective of our experimental evaluation is to answer three research questions:
@@
RQ2. How does FourierCSP compare to SAT, ILP, and CP solvers in solving constraint satisfaction problems that the constraints have structural patterns?
RQ3. How does FourierCSP perform relative to SAT, ILP, and CP solvers in solving constraint optimization problems that the constraints have structural patterns?
@@
To fully exploit the potential of Fourier expansion-based CLS, benchmark instances should ideally be symmetric, i.e., their constraints should exhibit similar structural patterns. This symmetry enables efficient GPU acceleration by allowing the solver to fully utilize parallelism, achieving up to 100+ times speedup in gradient computation compared to CPU (Cen, Zhang, and Fong 2025). The benchmarks include constraint satisfaction and optimization problems. The first benchmark is task scheduling problems, which can be useful for the trace scheduler of multicore hardware systems. For each combination of $N_{\text{worker}}\in\{32,64,128,256,512\}$ and $m\in\{4,8,16\}$, we set the number of jobs to $N_{\text{job}}=m\cdot N_{\text{worker}}$, and generate 10 random instances accordingly. The second benchmark, graph coloring with hashing queries, is an optimization problem motivated by approximate model counting (Pote, Meel, and Yang 2025), where hash functions are employed to partition the solution space. Specifically, we use graph coloring as the underlying problem and add parity (modulo 2) constraints to the formula to serve as hash functions. For each $N_{\text{node}}\in\{512,1024,2048,4096\}$ and $N_{\text{color}}\in\{8,16,32,64\}$, we generate 10 random instances. Additionally, $N_{\text{node}}$ modulo 2 constraints are added to each instance.
Benchmark instances are linearized for ILP solvers and Booleanized for Boolean SAT solvers. Further details on the benchmarks and the encodings used for different solvers can be found in the appendix, where we show that the Booleanization will naturally lead to pseudo-Boolean and XOR constraints. In particular, further encoding the largest Booleanized instance into CNF (Ignatiev, Morgado, and Marques-Silva 2018; Li 2000) yields a formula comprising over 160 million variables and 710 million clauses. Although modern SAT solvers are highly optimized and capable of solving large benchmarks, such encodings are extremely memory-inefficient; for instance, the corresponding CNF text file is around 16 GB in size.
@@
(1) FourierSAT (Kyrillidis et al. 2020): A CLS solver designed for Boolean formulas that are conjunctions of symmetric Boolean constraints. (2) FourierMaxSAT (Cen, Zhang, and Fong 2025): An extension of FourierSAT incorporating an SAT solver for warm-start initialization. (3) LinPB (Yang and Meel 2021): A SAT solver for linear pseudo-Boolean and XOR constraints. It employs RoundingSAT, which extends CDCL to pseudo-Boolean reasoning, as the underlying pseudo-Boolean solver. XOR constraints are supported via Gauss-Jordan elimination, adapted from CryptoMiniSAT. (4) Gurobi-ILP (Gurobi Optimization, LLC 2024): An ILP solver that uses branch-and-bound, cutting planes, and heuristics to handle integer linear constraints efficiently. The solver also leverages preprocessing and parallel computation to enhance performance. We use version 11.0.3. (5) OR-Tools CP-SAT (Perron, Didier, and Gay 2023): A hybrid solver that integrates a clause-learning SAT core, constraint programming, and linear relaxation within a unified model to solve combinatorial problems using advanced techniques; it won gold medals in all categories of the MiniZinc Challenge 2024. This solver won gold medals in all categories in the MiniZinc Challenge 2024. We use version 9.11. (6) ParaILP (Lin et al. 2024): A parallel discrete local search solver for ILP.
@@
Answer to RQ1.
Since FourierCSP is the multi-valued extension of FourierSAT, the semantics of solving hybrid SAT formulas remain equivalent to those of FourierSAT. Prior research indicates that FourierSAT demonstrates strong performance on benchmarks natively formulated as hybrid SAT problems. However, our experiments reveal a significant limitation: FourierSAT fails to solve any instances encoded from both scheduling and graph coloring benchmarks. This observation underscores a critical gap in the capability of FourierSAT when handling problems not inherently structured as conjunctions of symmetric Boolean constraints. Consequently, it highlights the necessity of extending the solver to a more general domain. Specifically, the results motivate the development of FourierCSP, a multi-valued extension of FourierSAT, which is designed to address precisely these limitations by supporting more general, i.e.. multi-valued formulations.
Figure 2: Results on graph coloring with hashing query problems. The numbers in the legend are (the average relative score, #Win of the solvers).
Answer to RQ2.
We assess solver performance on constraint satisfaction problems using the PAR-2 (penalized average runtime 2) score, a metric balancing runtime efficiency and robustness by imposing penalties on timeouts.
@@
Figure 1 presents the results on the first benchmark, constraint satisfaction problems. Specifically, FourierCSP and ParaILP represent two parallel local search solvers based on continuous and discrete optimization methods, respectively. Notably, under a timeout limit of 1000 seconds, FourierCSP consistently solves all 150 instances, substantially surpassing ParaILP, which solves only 60. This significant performance gap can be attributed to the differences in the search methods. Discrete local search methods require that each assignment strictly increase the number of satisfied constraints, often leading to plateaus, particularly for constraints that are long or structurally complex, where small perturbations to discrete variable assignments may not immediately yield progress. In contrast, CLS methods can exploit fine-grained improvements: the optimizer advances as long as the continuous objective function increases, even marginally.
@@
Answer to RQ3.
We evaluate this benchmark using two metrics: the relative score and the number of instances for which a solver attains the best objective value (#Win). The relative score is defined as follows.
$\text{score}(s,i)=\frac{1+c_{i}^{s}}{1+c_{i}^{*}}$
Notably, this benchmark includes a substantial number of modulo constraints, which are known to be challenging for CP-SAT (Krupke et al. 2024). While modulo constraints remain beyond the expressive or algorithmic reach of most conventional solvers, FourierCSP can natively accommodate them through its uniform Fourier expansion framework. As in Figure 2, FourierCSP outperforms all competing methods on this benchmark, achieving a relative score of 0.98 and attaining the best solution for 84 out of 160 instances In contrast to its competitive performance on the previous benchmark, CP-SAT’s relative score is 0.18 lower than that of Gurobi-ILP for these instances. This relative underperformance can likely be attributed to CP-SAT’s inefficiency in processing modulo constraints.
Under binary logarithm encoding, a modulo 2 constraint can be easily Booleanized to an XOR constraint. Although LinPB can natively solve XOR constraints using Gaussian elimination, it achieves a relative score of 0.16 on this benchmark. Nevertheless, LinPB contributes more to the virtual best solver than CP-SAT, as it attains the best solution for 17 instances. Notably, formulas with parity constraints are known to be challenging for DLS solvers (Crawford, Kearns, and Schapire 1994). It is unsurprising that ParaILP fails to provide valid solutions for any instance in this benchmark.
Previous work demonstrated that FourierSAT performs competitively on graph problems when the encoded formulas comprise exclusively symmetric Boolean constraints (Cen, Zhang, and Fong 2025). In contrast, the Booleanized formulas used in this benchmark consist primarily of pseudo-Boolean constraints, which would require further symmetrization into disjunctive clauses to fit the FourierSAT framework. This additional encoding step increases the complexity, making the problem significantly more challenging for FourierSAT, which does not provide valid solutions for any instance in this benchmark. While FourierSAT encounters difficulties when addressing formulas containing a large number of clauses, FourierMaxSAT adopts a hybrid approach. It leverages an SAT solver to solve part of the formula, i.e., the clauses, and subsequently uses CLS to improve the solution, i.e., by attempting to satisfy more XOR constraints. Eventually, FourierMaxSAT achieved a relative score of 0.94.
@@
Conclusion and Future Directions
In this paper, we introduced FourierCSP, a novel continuous optimization framework for constraint satisfaction problems. By extending the Walsh-Fourier expansion to finite-domain variables, our approach can handle versatile constraints without the need for memory-intensive encodings or auxiliary variables. The equivalence between circuit-output probability and Walsh-Fourier expansion allows us to evaluate and differentiate the objective function efficiently, enabling gradient-based optimization. We adopted a projected gradient approach for our optimization framework, which is supported by theoretical guarantees on convergence. Empirical evaluations on benchmark suites demonstrated that FourierCSP is scalable and performs competitively. The experimental results indicate that FourierCSP is a substantial extension of FourierSAT, broadening the class of problems that can be efficiently solved using CLS techniques.
The performance of Fourier transform-based CLS relies heavily on GPU parallelism. An important direction for future work is to investigate how this parallelism can be effectively leveraged when constraints exhibit non-structural patterns. This would further generalize FourierCSP, enabling its application to a broader range of CSP problems.

### v2

@@ 3.3 Optimization Approaches
$Q_{t+1}$ $=P_{t}-\eta\cdot\nabla\mathcal{F}(P_{t})$ (3)
$P_{t+1}$ $=\mathrm{proj}_{\tilde{D}}(Q_{t+1})$ (4)
Based on the randomized rounding in Definition 3.2, the Euclidean projection of $Q\in\mathbb{R}^{|\mathrm{dom}(x_{1})|\times\cdots\times|\mathrm{dom}(x_{n})|}$, onto $P\in\tilde{D}$, is defined as
@@
The projection operation $\mathrm{proj}_{\tilde{D}}(Q)$ can be computed efficiently by separately projecting each $q_{i}$ onto its respective simplex $\tilde{D}_{i}$, leveraging the product structure of $\tilde{D}$. Each sub-projection corresponds to a standard simplex projection problem (Chen and Ye, 2011).
Theorem 3.9.
@@
With the step size $\eta=\frac{1}{L}$, the PGD will converge to a $\epsilon$-critical point in $\mathcal{O}\left(\frac{|C|\cdot L}{\epsilon^{2}}\right)$ iterations.
Mirror Descent
While projected gradient methods rely on Euclidean geometry for projection, MD generalizes this process to non-Euclidean geometries induced by a strictly convex potential function $\psi$.
$Q_{t+1}$ $=\nabla\psi(P_{t})-\eta\nabla\mathcal{F}(P_{t}),$
$P_{t+1}$ $=\nabla\psi^{-1}(Q_{t+1}).$
We choose negative entropy as the potential function for each probability simplex $\tilde{D}_{i}$, hence $\psi(P)={\langle P,\log(P)\rangle}$. Then we have
$\nabla\psi(P)=\log(P),\quad\nabla\psi^{-1}(Q)=\frac{\exp(Q)}{\langle\exp(Q),\mathbf{1}\rangle}.$
Theorem 3.10.
Let $\Theta=\sum_{i=1}^{n}\log|\mathrm{dom}(x_{i})|$. Let $T$ be the number of MD steps. with the step size $\eta\leq\frac{\sqrt{2\Theta}}{\rho\sqrt{T}}$, then MD will converge to a $\epsilon$-critical point in $\mathcal{O}\left(\frac{2\Theta\cdot\rho}{\epsilon^{2}}\right)$ iterations.
Hybrid Descent
Theorem 3.9 implies a larger gradient norm $\|\nabla\mathcal{F}\|$ leads to a greater decrease in the objective at each projected gradient step, and therefore accelerates convergence. In contrast, Theorem 3.10 requires a global bound $\|\nabla\mathcal{F}\|\leq\rho$, and a larger gradient increases this bound, which forces a smaller step-size in the analysis and ultimately slows the convergence rate of MD. This contrast is made explicit in (Allen-Zhu and Orecchia, 2014), which shows that PGD and MD operate under different geometries and therefore complement each other. Motivated by this observation, we adopt a hybrid update rule, HD, that evaluates both descent directions and selects the better one:
$P_{PGD}=\mathrm{proj}_{\tilde{D}}(P_{t}-\eta_{PGD}\cdot\nabla\mathcal{F}(P_{t})),$
$P_{MD}=\nabla\psi^{-1}(\nabla\psi(P_{t})-\eta_{MD}\nabla\mathcal{F}(P_{t})),$
$P_{t+1}=\operatorname*{arg\,min}_{P\in\{\,P_{\mathrm{PGD}},\,P_{\mathrm{MD}}\}}\mathcal{F}(P).$
Implementation Details
FourierCSP is implemented by extending the core jaxopt.ProjectedGradient framework and incorporating the projection and mapping functions from jaxopt.MirrorDescent (Bradbury et al., 2018; Blondel et al., 2022). We adopt the default settings in jaxopt, using a maximum of 500 gradient steps and a tolerance of 0.001. The step size is determined automatically via backtracking line search, with a maximum of 15 backtracking steps and a decrease factor of 0.5. All descent methods are evaluated both with and without Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) acceleration.
4 Experimental Section
@@
In this section, we compare our tool, FourierCSP, with versatile state-of-the-art solvers. The objective of our experiments is to answer the following research questions.
@@
RQ2. How do different descent methods compare in effectiveness and robustness in the FourierCSP framework?
RQ3. How does FourierCSP compare to SAT, ILP, and CP solvers in solving constraint satisfaction problems that the constraints have structural patterns?
RQ4. How does FourierCSP perform relative to SAT, ILP, and CP solvers in solving constraint optimization problems that the constraints have structural patterns?
(a) #Iter to local minimum
(b) Time to local minimum
(c) #Iter to global minimum
(d) Time to global minimum
Figure 1: The performance of FourierCSP with different descent methods on the random benchmark.
@@
To fully exploit the potential of Fourier expansion-based CLS, benchmark instances should ideally be symmetric, i.e., their constraints should exhibit similar structural patterns. This symmetry enables efficient GPU acceleration by allowing the solver to fully utilize parallelism, achieving up to 100+ times speedup in gradient computation compared to CPU (Cen et al., 2025). The benchmarks include constraint satisfaction and optimization problems. The first benchmark consists of random hybrid constraints, combining greater-than, not-all-equal, and modulo constraints. For each combination of $n_{\text{var}}\in\{100,...,1000\}$ and $n_{\text{state}}\in\{4,8,16,32\}$, we generate 10 random instances. Each instance contains $\frac{n_{\text{var}}}{10}$ of each type. The arity (length) of NAE and modulo constraints is fixed to 10. The second benchmark is task scheduling problems, which can be useful for the trace scheduler of multicore hardware systems. For each combination of $N_{\text{worker}}\in\{32,64,128,256,512\}$ and $m\in\{4,8,16\}$, we set the number of jobs to $N_{\text{job}}=m\cdot N_{\text{worker}}$, and generate 10 random instances accordingly. The third benchmark, graph coloring with hashing queries, is an optimization problem motivated by approximate model counting (Pote et al., 2025), where hash functions are employed to partition the solution space. We use graph coloring as the underlying problem and add parity (modulo 2) constraints to the formula to serve as hash functions. For each $N_{\text{node}}\in\{512,1024,2048,4096\}$ and $N_{\text{color}}\in\{8,16,32,64\}$, we generate 10 random instances. Additionally, $N_{\text{node}}$ modulo 2 constraints are added to each instance.
Benchmark instances are linearized for ILP solvers and Booleanized for Boolean SAT solvers. Further details on the benchmarks and the encodings used for different solvers can be found in the appendix, where we show that the Booleanization will naturally lead to pseudo-Boolean and XOR constraints. Further encoding the largest Booleanized instance into CNF (Ignatiev et al., 2018; Li, 2000) yields a formula comprising over 160 million variables and 710 million clauses. Although modern SAT solvers are highly optimized and capable of solving large benchmarks, such encodings are extremely memory-inefficient; for instance, the corresponding CNF text file is around 16 GB in size. The details of the encoding are in Appendix B.
@@
(1) FourierSAT (Kyrillidis et al., 2020): A CLS solver designed for Boolean formulas that are conjunctions of symmetric Boolean constraints. In (Cen et al., 2025), it uses PGD with FISTA as the optimizer. (2) FourierMaxSAT (Cen et al., 2025): An extension of FourierSAT incorporating an SAT solver for warm-start initialization. (3) LinPB (Yang and Meel, 2021): A SAT solver for linear pseudo-Boolean and XOR constraints. It employs RoundingSAT, which extends CDCL to pseudo-Boolean reasoning, as the underlying pseudo-Boolean solver. XOR constraints are supported via Gauss-Jordan elimination, adapted from CryptoMiniSAT. (4) Gurobi-ILP (Gurobi Optimization, LLC, 2024): An ILP solver that uses branch-and-bound, cutting planes, and heuristics to handle integer linear constraints efficiently. The solver also leverages preprocessing and parallel computation to enhance performance. We use version 11.0.3. (5) OR-Tools CP-SAT (Perron et al., 2023): A hybrid solver that integrates a clause-learning SAT core, constraint programming, and linear relaxation within a unified model to solve combinatorial problems using advanced techniques; it won gold medals in all categories of the MiniZinc Challenge 2024. This solver won gold medals in all categories in the MiniZinc Challenge 2024. We use version 9.11. (6) ParaILP (Lin et al., 2024): A parallel discrete local search solver for ILP.
@@
We assess solver performance on satisfaction problems using the PAR-2 (penalized average runtime 2) score, a metric balancing runtime efficiency and robustness by imposing penalties on timeouts.
@@
We evaluate the optimization problem using two metrics: the relative score and the number of instances for which a solver attains the best objective value (#Win). The relative score is defined as follows.
$\text{score}(s,i)=\frac{1+c_{i}^{s}}{1+c_{i}^{*}}$
(a) FourierCSP with different descent methods.
(b) Comparison of FourierCSP with other solvers.
4.1 Answer to RQ1.
Since FourierCSP is the multi-valued extension of FourierSAT, the semantics of solving hybrid SAT formulas remain equivalent to those of FourierSAT. Prior research indicates that FourierSAT demonstrates strong performance on benchmarks natively formulated as hybrid SAT problems. However, our experiments reveal a significant limitation: FourierSAT fails to solve any instances encoded from both scheduling (Fig. 2) and graph coloring benchmarks (Fig. 3). This observation underscores a critical gap in the capability of FourierSAT when handling problems not inherently structured as conjunctions of symmetric Boolean constraints. Consequently, it highlights the necessity of extending the solver to a more general domain. Specifically, the results motivate the development of FourierCSP, a multi-valued extension of FourierSAT, which is designed to address these limitations precisely by supporting more general, i.e.. multi-valued formulations.
4.2 Answer to RQ2.
From Fig. 1 (a), MD requires the fewest gradient iterations to reach a local minimum. However, Fig. 1 (b) shows that it nevertheless takes longer in wall-clock time. This is because all methods use line search to select the step size: although MD makes rapid progress per gradient step, its updates satisfy the line-search conditions less easily, causing more backtracking. As a result, each MD iteration is more expensive, leading to a higher total runtime despite fewer iterations. It’s counterintuitive that HD is the most time-efficient approach, where each gradient step needs to measure the gradient step made by both PGD and MD. This is because HD inherits the faster descent direction of the two while avoiding the costly backtracking behavior that slows down pure MD. The same pattern appears in Fig. 1(c–d) for convergence to the global minimum.
Across all settings, HD is the most effective and robust method within the FourierCSP framework for constraint satisfaction problems, consistently solving the largest number of instances under a fixed trial budget. FISTA further accelerates both PGD and MD, with PGD-fista performing competitively with HD and HD-fista. This trend is confirmed on the scheduling benchmark in Fig. 2 (a), where plain PGD is significantly slower than all other descent methods, while PGD-fista eliminates most of this gap, and even achieves the lowest PAR-2 score.
The results for the constraint optimization benchmark are reported in Table 1. Although PGD-fista achieves the highest anytime score, MD-fista yields the largest number of best solutions. Moreover, other descent methods also achieve the best result on certain instances, indicating that no single method uniformly dominates across all problems. These observations suggest that a portfolio strategy combining multiple descent methods is more reliable and robust than relying on any individual optimizer.
[…]
Descent Method Anytime Score #Win
PGD 0.87 26
PGD-fista 0.97 15
MD 0.80 20
MD-fista 0.76 46
HD 0.86 11
HD-fista 0.92 13
Table 1: Performance of FourierCSP with different descent methods on the graph coloring with hashing query benchmark.
Figure 3: Results on graph coloring with hashing query problems. The numbers in the legend are (the average relative score, #Win) of the solvers.
4.3 Answer to RQ3.
Fig. 2 presents the results on the second benchmark, constraint satisfaction problems. FourierCSP and ParaILP represent two parallel local search solvers based on continuous and discrete optimization methods, respectively. Under a timeout limit of 1000 seconds, FourierCSP successfully solves all 150 instances, substantially surpassing ParaILP, which solves only 60. This significant performance gap can be attributed to the differences in the local search methods. Discrete local search methods require that each assignment strictly increase the number of satisfied constraints, often leading to plateaus, particularly for constraints that are long or structurally complex, where small perturbations to discrete variable assignments may not immediately yield progress. In contrast, CLS methods can exploit fine-grained improvements: the optimizer advances as long as the continuous objective decreases, even marginally.
@@
4.4 Answer to RQ4.
This benchmark includes a substantial number of modulo constraints, which are known to be challenging for OR-Tools CP-SAT (Krupke et al., 2024). While modulo constraints remain beyond the expressive or algorithmic reach of most conventional solvers, FourierCSP can natively accommodate them through its uniform Fourier expansion framework. As shown in Fig. 3, FourierCSP-pf outperforms all competing solvers on this benchmark, achieving a relative score of 0.96. The FourierCSP-vbs setting reflects the performance upper bound of FourierCSP-pf and attains the best solution on 120 out of 160 instances. In contrast to its competitive performance on the previous benchmark, OR-Tools CP-SAT’s relative score is 0.18 lower than that of Gurobi-ILP for these instances. This relative underperformance can likely be attributed to OR-Tools CP-SAT’s inefficiency in processing modulo constraints.
Under binary logarithm encoding, a modulo 2 constraint can be easily Booleanized to an XOR constraint. Although LinPB can natively solve XOR constraints using Gaussian elimination, it achieves a relative score of 0.16 on this benchmark. Nevertheless, LinPB contributes more to the virtual best solver than OR-Tools CP-SAT, as it attains the best solution for 14 instances. Notably, formulas with parity constraints are known to be challenging for DLS solvers (Crawford et al., 1994). Unsurprisingly, ParaILP fails to provide valid solutions for any instance in this benchmark.
Previous work demonstrated that FourierSAT performs competitively on graph problems when the encoded formulas comprise exclusively symmetric Boolean constraints (Cen et al., 2025). In contrast, the Booleanized formulas used in this benchmark consist primarily of pseudo-Boolean constraints, which would require further symmetrization into disjunctive clauses to fit the FourierSAT framework. This additional encoding step increases the complexity, making the problem significantly more challenging for FourierSAT, which does not provide valid solutions for any instance in this benchmark. While FourierSAT encounters difficulties when addressing formulas containing a large number of clauses, FourierMaxSAT adopts a hybrid approach. It leverages an SAT solver to solve part of the formula, i.e., the disjunctive clauses, and subsequently uses CLS to improve the solution, i.e., by attempting to satisfy more XOR constraints. Eventually, FourierMaxSAT achieved a relative score of 0.93.
@@
5 Conclusion and Future Directions
In this paper, we introduced FourierCSP, a novel continuous optimization framework for constraint satisfaction problems. By extending the Walsh-Fourier expansion to finite-domain variables, our approach can handle versatile constraints without the need for memory-intensive encodings or auxiliary variables. By randomized rounding, the discrete constraints are transformed into differentiable functions, enabling gradient-based optimization. We adopted multiple gradient-based approaches for our optimization framework, which are supported by theoretical guarantees on convergence. Empirical evaluations on benchmark suites demonstrated that FourierCSP is scalable and performs competitively. The experimental results indicate that FourierCSP is a substantial extension of FourierSAT, broadening the class of problems that can be efficiently solved using CLS techniques. Ultimately, the differentiable nature of FourierCSP facilitates seamless integration into neurosymbolic architectures, bridging the long-standing gap between data-driven deep learning and constraint reasoning.
The performance of Fourier transform-based CLS relies heavily on GPU parallelism. An important direction for future work is to investigate how this parallelism can be effectively leveraged when constraints exhibit non-structural patterns. One promising approach is to leverage compact symbolic data structures, such as decision diagrams, to represent constraints and compute their circuit-output probabilities efficiently (Appendix C). This would further generalize FourierCSP, enabling its application to a broader range of CSP problems.